# Error Correction Code Transformer: From Non-Unified to Unified

### v1

@@ 1.1 Prior Works
Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms. Looking ahead to 6G, this trend is likely to persist [8]. For instance, the Successive Cancellation (SC) decoding algorithm [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result. This approach simplifies the hardware implementation of Polar codes by pruning binary trees [9]. In contrast, the Sum-Product (SP) decoding algorithm [10] for LDPC codes uses iterative message passing, allowing for parallel data processing. This fundamentally differs from the sequential method used for Polar codes in hardware implementation [11]. Furthermore, BCH codes typically use algebraic decoding techniques [12], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP. Overall, these algorithms provide strong error correction capabilities. However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code. This may lead to significant engineering effort and inefficient use of hardware resources.
In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, noise estimation, and Artificial Intelligence (AI)-driven approaches. Resource sharing involves time-sharing storage and computing units. Noise estimation methods decode various linear block codes by estimating the noise of received signals. Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
@@
Second, noise estimation methods were first proposed in [15] as the Guessing Random Additive Noise Decoding (GRAND) algorithm, used for Maximum Likelihood (ML) decoding of arbitrary codebooks in discrete channels, with or without memory [16, 17]. The algorithm operates by rank-ordering noise sequences based on their likelihood and subtracting them from the received signal to identify codebook members [16]. GRAND is shown to be capacity-achieving with random codebooks. However, this work faces the challenge of excessive guessing required by the GRAND algorithm, particularly for high logistic weights, which can lead to computational inefficiency. For instance, when the logistic weight equals $128$, up to $5.33\times 10^{7}$ queries may be necessary to complete a single decoding process [17].
Third, in recent years, AI-driven decoders have gained attention as an alternative. While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models [18], [19]. In contrast, model-free approaches offer a more flexible and scalable solution. Unlike model-based methods, which depend on predefined mathematical models, model-free approaches do not assume prior knowledge of the encoder or decoder. Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms [20]. For example, inspired by the success of Transformers in Natural Language Processing (NLP) [21], a Transformer-based, model-free neural decoder called the Error Correction Code Transformer (ECCT) is proposed [22]. ECCT leverages the Transformer’s self-attention mechanism, where multiple attention heads work in parallel to effectively manage sequential data and capture long-range dependencies, to establish relationships between the input Log-Likelihood Ratios (LLRs) in the decoder. The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits. However, despite offering greater flexibility, model-free approaches like ECCT still require frequent tuning of model parameters for each code type, length, and rate. This increases training costs and hardware overhead, making the decoding architecture non-unified.
@@ 1.2 Contributions
    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture. Additionally, we eliminated the need for iterations through the unified attention mechanism, thereby avoiding the detrimental effects of short cycles in Tanner graphs on LDPC belief propagation decoders. As a result, we have unlocked the full potential of LDPC codes, extending their applicability to short code scenarios.
@@
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $83\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
@@
Notations: Lowercase and uppercase bold letters represent vectors and matrices, respectively. GF(2) denotes the binary Galois field. $\mathbb{R}$ represents real numbers. $[\cdot]^{T}$ stands for transpose, while $\mathcal{N}(\mu,{\sigma}^{2})$ signifies a Gaussian distribution with mean $\mu$ and variance ${\sigma}^{2}$. We define $bipolar(\cdot)$ to map binary to bipolar signals, and $bin(\cdot)$ represents its inverse mapping, where $bipolar(x)$ is defined as $1-2x$, and $x$ takes values in the set $\{0,1\}$. $size(\cdot)$ represents the number of elements in a matrix or vector. The sign function is denoted as $sign(\cdot)$.
@@ 2.1 Vanilla Transformers
The vanilla Transformer functions as a sequence-to-sequence model comprising an encoder and a decoder, both built from a series of $L$ identical blocks. Each encoder block primarily integrates a multi-head self-attention module and a position-wise feed-forward network (FFN). To construct a deeper model and prevent gradient degradation, residual connections[23] are employed between the self-attention module and the FFN. Layer normalization[24] is applied to the inputs of each module, enhancing network stability and convergence speed. The overall architecture of the single-layer vanilla Transformer encoder is illustrated in Figure 1.
Figure 1: Overall architecture of the single-layer vanilla Transformer.
The following subsections will introduce the key modules within the vanilla Transformer.
2.1.1 Word Embedding
Word embedding is an important term in natural language processing, used to represent words for text analysis in the form of real-valued vectors. It enhances the computer’s ability to better understand text-based content. In this paper, word embeddings serve as reliability encoders that map the received one-dimensional signal to higher dimensions. It can be expressed as
$Embedding\left(\textbf{{X}}\right)=\textbf{{X}}\odot\textbf{{E}}$ (1)
where $\textbf{{X}}\in\mathbb{R}^{N\times 1}$ is the received signal with length $N$ and dimension one, $\textbf{{E}}\in\mathbb{R}^{N\times d_{k}}$ represents the trainable embedding matrix with dimensions of $d_{k}$, and $\odot$ denotes the Hadamard product.
2.1.2 Multi-Head Self-Attention
The vanilla Transformer establishes an attention mechanism based on the three Query-Key-Value (Q-K-V) matrices. It utilizes the Q and K matrices to establish correlations between different elements within the sequence and weights the V matrix accordingly. The scaled dot-product form of attention is given by
$Attention\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Softmax\left(\frac{\textbf{{QK}}^{T}}{\sqrt{d_{k}}}\right)\textbf{{V}}$ (2)
$Softmax(\textbf{{z}})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{n}{e^{z_{j}}}}~{}for~{}i=1,2,\cdots,n$ (3)
where $\textbf{{Q}}\in\mathbb{R}^{N\times d_{k}}$, $\textbf{{K}}\in\mathbb{R}^{M\times d_{k}}$, $\textbf{{V}}\in\mathbb{R}^{M\times d_{v}}$ represent tensors with lengths of $N$ (for queries and keys) and $M$ (for values), and dimensions of $d_{k}$ (for queries and keys) and $d_{v}$ (for values), respectively. The $Softmax\left(\cdot\right)$ function normalizes inputs into the range of 0 to 1, and to alleviate its gradient degradation issue, the dot product of queries and keys is divided by $\sqrt{d_{k}}$.
The study in[21] enhanced the attention layer by introducing a technique known as multi-head attention, which involves $H$ different learned projection sets. This enhancement benefits the attention layer’s performance in two significant ways: it widens the model’s capacity to concentrate on various positions and grants the attention layer multiple representation subspaces. The multi-head attention can be written as
$MultiHeadAttn\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Concat\left(head_{1},head_{2},\cdots,head_{H}\right)\textbf{{W}}^{\textit{O}}$ (4)
where $head_{i}=Attention(\textbf{{QW}}_{i}^{\textit{Q}},\textbf{{KW}}_{i}^{\textit{K}},\textbf{{VW}}_{i}^{\textit{V}})$, and $\textbf{{W}}_{i}^{\textit{Q}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{K}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{V}}\in\mathbb{R}^{d_{v}\times d_{v}}$ are trainable parameters, $\textbf{{W}}^{\textit{O}}\in\mathbb{R}^{Hd_{v}\times Hd_{v}}$ is a weight matrix that was trained jointly with the model, serving to condense the outputs of different heads down into a single matrix.
@@ 2.1.2 Multi-Head Self-Attention
2.1.3 Position-Wise Feed-Forward Network
The outputs of the self-attention layer are fed into the position-wise FFN layer. The position-wise FFN layer is essentially a fully connected network that operates uniformly on each position, and the non-linear activation function within it enables the Transformer to learn more complex feature representations. The mathematical expression for the FFN is as follows
$FFN\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\times ReLU\left(\textbf{{W}}_{1}\times\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (5)
$ReLU\left(x\right)=max\left(0,x\right)$ (6)
where X is the output of previous self-attention layer, and $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}\in\mathbb{R}^{d_{f}}$, $\textbf{{b}}_{2}\in\mathbb{R}^{d_{h}}$ are trainable parameters, $d_{h}=H\times d_{k}$ is the dimensions of queries, keys and values, $ReLU\left(\cdot\right)$ represents the activation function. In general, to achieve better performance, $d_{f}$ is set to be greater than $d_{h}$.
2.1.4 Residual Connection and Normalization
The multi-head self-attention layer and the position-wise FFN layer are connected through residual network[23] and layer normalization[24]. The overall architecture of each vanilla Transformer encoder can be written as
$\textbf{{X}}_{1}$ $=LayerNorm\left(MultiHeadAttn(\textbf{{X}})+\textbf{{X}}\right)$ (7)
$\textbf{{X}}_{2}$ $=LayerNorm\left(FFN(\textbf{{X}}_{1})+\textbf{{X}}_{1}\right)$
where $LayerNorm\left(\cdot\right)$ denotes the layer normalization module.
2.2 Forward Error Correction Codes
In information theory, FEC is a technique designed to reduce the bit error rate (BER) when transmitting data over a noisy channel. We establish a standard linear block code utilizing a generator matrix G with dimensions $k\times n$ and a parity-check matrix H with dimensions $(n-k)\times n$. These matrices are specifically designed to satisfy $\textbf{G}\textbf{H}^{T}=\textbf{0}$ over the binary Galois field (GF).
We will now briefly discuss the encoder. The message vector $\textbf{{m}}\in\{0,1\}^{k}$ multiplied with the generator matrix G over the GF(2) results in a codeword vector $\textbf{{x}}\in\{0,1\}^{n}$ satisfying $\textbf{H}\textbf{{x}}=\textbf{0}$. The encoded codeword x is modulated by Binary Phase Shift Keying (BPSK) to $\textbf{{x}}_{s}$, where $BPSK(x)=1-2x$. $\textbf{{x}}_{s}$ is transmitted through binary-input discrete memoryless channels (B-DMCs, e.g. an Additive White Gaussian Noise channel), resulting in a channel output $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, where n is Gaussian noise with zero mean and variance $\sigma^{2}$.
@@ 2.2 Forward Error Correction Codes
The goal of the decoder is to recover the transmitted message m from the received signal y. The overall architecture of the communication system including both encoding and decoding is illustrated in Figure 3. The objective of our work is to design and training of the parameterized decoding function $\textit{f}_{\theta}$.
@@
To build a model-free Transformer decoding model, we draw inspiration from the pre- and post-processing methods proposed in [25], which have been demonstrated not to result in performance loss concerning bit error rate or mean squared error (MSE) metrics, as depicted in Figure 4. By training with an all-zero codeword to learn the noise, it is possible to avoid the overfitting problem that arises from the exponential growth of the codeword space with code length. And the B-DMCs is modeled by (8) which is an equivalent statistical mode that differs from the true physical one [26].
@@
In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as
$=\left[\lvert\textbf{{y}}\rvert,s\left(\textbf{{y}}\right)\right]$ y ) ] (9)
@@
In the post-processing step, the predicted multiplicative noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form
$\hat{\textbf{{x}}}=bin\left(sign\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (10)

### v2

@@ 1.1 Prior Works
Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms. Looking ahead to 6G, this trend is likely to persist [8]. For instance, the Successive Cancellation (SC) decoding algorithm [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result. This approach simplifies the hardware implementation of Polar codes by pruning binary trees [9, 10, 11]. In contrast, the Sum-Product (SP) decoding algorithm [12] for LDPC codes uses iterative message passing, allowing for parallel data processing. This fundamentally differs from the sequential method used for Polar codes in hardware implementation [13]. Furthermore, BCH codes typically use algebraic decoding techniques [14], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP. Overall, these algorithms provide strong error correction capabilities. However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code. This may lead to significant engineering effort and inefficient use of hardware resources.
In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, Ordered-Statistics Decoding (OSD), and Artificial Intelligence (AI)-driven approaches. Resource sharing involves time-sharing storage and computing units. Ordered-Statistics Decoding leverages the reliability ordering of received symbols to efficiently decode various linear block codes by systematically evaluating the most likely codewords. Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
@@
Second, ordered-statistics decoding methods, initially explored in [17, 18, 19], provide a near-Maximum Likelihood (ML) decoding approach for linear block codes by exploiting the reliability ordering of received symbols. The algorithm operates by sorting the received signal positions based on their reliability, permuting the code’s generator matrix accordingly, and systematically evaluating the most likely codewords through a limited search. OSD delivers near-ML performance with significantly less complexity than exhaustive ML decoding, making it well-suited for short to moderate-length codes. However, its drawback lies in the computational complexity, which grows exponentially with the search order (e.g., order-$l$). For instance, an order-4 OSD may require evaluating thousands of test codewords, potentially leading to increased latency and resource demands.
Third, in recent years, AI-driven decoders have gained attention as an alternative. While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models [20], [21]. In contrast, model-free approaches offer a more flexible and scalable solution. Unlike model-based methods, which depend on predefined mathematical models, model-free approaches assume no prior knowledge of the encoder or decoder. Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms [22].
For instance, inspired by the success of Transformers in Natural Language Processing (NLP) [23], a model-free neural decoder called the Error Correction Code Transformer (ECCT) was proposed [24]. ECCT employs the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively linking input Log-Likelihood Ratios (LLRs) to the decoder’s output. The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits. However, ECCT’s flexibility comes at a cost, requiring frequent parameter adjustments for different code types, lengths, and rates, which increases training complexity and hardware overhead, hindering a unified decoding architecture. Building on this, [25] introduced a Foundation Model for Error Correction Codes (FECCT) to improve ECCT’s generalization to unseen codewords. Yet, FECCT currently faces the challenge of high computational complexity in its self-attention mechanism, further complicating its practical implementation.
@@ 1.2 Contributions
    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture.
@@
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
@@
Notations: Bold lowercase and uppercase letters denote vectors (e.g., $\mathbf{v}$) and matrices (e.g., $\mathbf{M}$), respectively, while scalars use regular letters (e.g., $x$). $\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers. The transpose is indicated by $[\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values ${1,-1}$, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping ${1,-1}$ back to ${0,1}$. The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 00 if $x=0$, and $-1$ if $x<0$. For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
@@ 2.1 Vanilla Transformers
The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks. Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections [26] to mitigate gradient issues, and layer normalization [27] for stability and faster convergence, as shown in Figure 1.
Figure 1: Single-layer vanilla Transformer encoder architecture.
The process begins with embedding a one-dimensional input signal $\textbf{{x}}\in\mathbb{R}^{N}$ into a higher-dimensional space:
$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}}$ (1)
where $\text{diag}(\textbf{{x}})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with x on the diagonal, and $\textbf{{W}}\in\mathbb{R}^{N\times d_{k}}$ is a trainable weight matrix, producing $\textbf{{X}}_{e}\in\mathbb{R}^{N\times d_{k}}$ with $d_{k}$-dimensional embeddings for each position.
This embedding feeds into the MHA mechanism, which correlates sequence elements through:
$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$ (2)
extended to multiple heads as:
$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
where $\text{head}_{i}=\text{Attn}(\textbf{{QW}}_{i}^{Q},\textbf{{KW}}_{i}^{K},\textbf{{VW}}_{i}^{V})$, with $\textbf{{Q}},\textbf{{K}},\textbf{{V}}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
@@
Next, the MHA output is processed by the FFN, applied uniformly across positions:
$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (4)
where $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}$, $\textbf{{b}}_{2}$ are trainable parameters, $d_{h}=H\cdot d_{k}$ and $\text{ReLU}(x)=\max(0,x)$ enables complex feature learning, with $d_{f}>d_{h}$ for enhanced performance.
Finally, residual connections and layer normalization link the MHA and FFN layers:
$\textbf{{X}}_{1}$ $=\text{MHA}\left(\text{LayerNorm}\left(\textbf{{X}}_{e}\right)\right)+\textbf{{X}}_{e}$ (5)
$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}$
2.2 Forward Error Correction Codes
In information theory, FEC is a technique designed to reduce the bit error rate (BER) when transmitting data over a noisy channel. We adopt a linear block code defined by a generator matrix $\textbf{G}\in\mathbb{R}^{k\times n}$ and a parity-check matrix $\textbf{H}\in\mathbb{R}^{(n-k)\times n}$, satisfying $\textbf{G}\cdot\textbf{H}^{T}=\textbf{0}$ over $\mathbb{F}_{2}$.
@@ 2.2 Forward Error Correction Codes
The decoder aims to recover m from y, as depicted in Figure 3. Our work focuses on designing and training a parameterized decoding function $f_{\theta}$. To develop a model-free Transformer-based decoder, we leverage pre- and post-processing techniques from [28], preserving BER and Mean Squared Error (MSE) performance (see Figure 4). Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth. And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one [29].
@@
In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as:
y~=[|y|,s($\textbf{{y}}\right)\right]$ = [ | y | , s ( y ) ] (7)
@@
In the post-processing step, the predicted noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form:
$\hat{\textbf{{x}}}=\text{bin}\left(\text{sign}\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (8)
2.3 Mask Generation in ECCT and FECCT
Masks in Transformer-based error correction enhance decoding by integrating code structure into attention mechanisms. In ECCT [24], the mask $g(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{(2n-k)\times(2n-k)}$ embeds code structure using the parity-check matrix H. Initialized as an identity matrix, it unmasks positions of paired ones in H rows, extending the Tanner graph to two-ring connectivity. It is applied additively in attention: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}+g(\textbf{H})}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$.
In FECCT [25], a learned $\psi:\mathbb{N}\to\mathbb{R}$ modulates attention via the distance matrix $\mathcal{G}(\textbf{H})\in\mathbb{N}^{(2n-k)\times(2n-k)}$, applied with a Hadamard product: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\left(\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\odot\psi(\mathcal{G}(\textbf{H}))\right)\cdot\textbf{{V}}$.